OpenStudy (anonymous):

anyone know how to this one :( Find the intercepts, and then graph the line. -3x-9y=9

OpenStudy (anonymous):

Simplify the equation and solve for y. -3x-9y=9 -9y=-3x+9 y=3/9x-9/9 y=1/3x-1 The equation of the line is now in slope intercept form and y=1/3x-1.

OpenStudy (anonymous):

kk.

OpenStudy (anonymous):

The slope is rise over run, and in this case our slope equals 1/3, so the line travels up 1 and over 3. The y-intercept is (0, -1). When you graph the line, start at (0, -1) and go up 1 and over 3 and make a point. Keep doing this, then connect them with a straight line.

OpenStudy (anonymous):

k